Ann has 120 roses, while Emily has 160 roses. How many roses should Emily give to Ann so that Emily will have 3 times less roses than Ann will have?
Answer by Boreal(15235)

gives her x roses and assume 3 times fewer means 1/3 as many.
then (160-x)/(120+x)=1/3
cross-multiply
480-3x=120+x
4x=360
x=90 ANSWER
then Emily will have 70 and Ann will have 210